OpenStudy (anonymous):

Explain how to factor the following trinomial forms: x2 + bx + c. Is there more than one way to factor this?

jimthompson5910 (jim_thompson5910):

Method 1: Find two numbers that multiply to c AND also add to b call these two numbers p and q ... ie p*q = c and p+q = b this would mean that x^2 + bx + c factors to (x+p)(x+q)

jimthompson5910 (jim_thompson5910):

Method 2: Find the x-intercepts of y = x^2 + bx + c by graphing if the x-intercepts are p and q, then x^2 + bx + c factors to (x-p)(x-q)

jimthompson5910 (jim_thompson5910):

Method 3: Use the quadratic formula to find the two solutions x = p or x = q, which means x^2 + bx + c factors to (x-p)(x-q)

OpenStudy (anonymous):

Show your answer using both words and mathematical notation @jim_thompson5910